# Hyper-Compression: Model Compression via Hyperfunction

### v1

@@ I. Limit of the Compression Ratio of Pruning Convolutional Networks
The class of graphs we investigate is $G_{r}=(V_{r},E_{r}),~{}r\in\mathbb{Z}_{\geq 2}$, where $V_{r}=\mathbb{Z}^{2},E_{r}=\{\langle(m,n),(m+1,n+i)\rangle\mid m,n\in\mathbb{Z},i\in{0,1,\cdots,r-1}\}$, that is, a convolution with the kernel size $r$ and stride 1.
@@
Earlier, researchers have mainly focused on planar graphs, but we observe that $G_{r}$ is a planar graph only when $r=2$, and is non-planar for $r\geq 3$, part of $G_{2}$ and $G_{3}$ are shown in Figures 5 and 5. Therefore, many methods, such as using the duality of planar graphs and the star-triangular transition, cannot be applied. In the following, we will first solve the critical probability for the case of $r=2$, and then study the upper and lower bounds of the critical probability for $r\geq 3$. Our method is degenerating a non-planar graph into a planar graph by removing edges.
@@
Remark. From the viewpoint of percolation theory, the compression ratio of pruning a randomly generated convolutional network is $1/p_{0}<\frac{1}{p_{c,r}}<\mu_{r}$, which is a fixed constant given $r$. Our analysis is based on an assumption that ignores the association between two edges. In reality, when an edge with a larger weight is pruned, edges with smaller weights are also pruned; otherwise, they are not. Future work will consider the association to have a more accurate estimation.
II. The Trend of Growth of Large Models
In recent years, the field of natural language processing (NLP) has witnessed remarkable advancements, particularly with the development of large language models (LLMs). These large models have demonstrated impressive capabilities in various language-related tasks, with a rapid increase in size over the past few years. For instance, GPT-3 consists of a staggering 175 billion parameters. Recently, M6, one of the largest models to date, has over $10^{4}$ billion parameters. This growth in model size has been driven by the desire to improve the model’s complex language understanding by scaling law. However, the growth of GPU memory has not kept pace. GPUs, which are crucial for training and serving these models, have seen relatively modest increases in memory capacity. This disparity between models’ size and GPU memory has created significant challenges in the development of large models.
Specifically, the disparity between the model size and GPU memory brings the following problems:
  1. 1.
    Memory Bottleneck: The limited GPU memory poses a bottleneck for serving LLMs. When deploying LLMs for real-world applications, the entire model and its associated parameters need to fit within the GPU memory. However, as LLMs continue to grow in size, it becomes increasingly difficult to accommodate them within the available memory constraints. This limitation hampers the deployment and wide adoption of LLMs in practical scenarios.
  2. 2.
    Ethic Issues: Since large models are hard to serve, large institutes will gradually dominate the development and deployment of large language models, which raises ethical concerns:
    Algorithmic Bias: LLMs developed by dominant institutes may reflect the biases present in the data they were trained on. This can perpetuate existing societal biases and discrimination, leading to unfair outcomes in decision-making processes.
    Lack of Transparency: Large institutes may not always disclose the full details of their model architectures, training data, or decision-making processes. This lack of transparency can hinder accountability, making it difficult to assess the fairness and reliability of LLMs in real-world applications.
As the demand for more powerful models and the risk of monopoly by large institutes continue to rise, addressing the disparity between model size and GPU memory is a paramount research topic. Particularly, we need to cultivate novel model compression techniques that can realize Moore’s law of model compression, facilitating the deployment and efficient utilization of LLMs in practical scenarios. Moore’s law of model compression, which proposes that the compression ratios of models should double annually, is the potential solution. By compressing models, we can reduce their size and memory requirements, making them more feasible to deploy on GPUs with limited memory.
@@ III. Algorithmic Details
Anchored in the hyperfunction driven by ergodic theory mentioned in the main body, we design a general-purpose model compression method that enjoys the following merits (PNAS): Preferable compression ratio, No post-hoc retraining, Affordable inference time, and Short compression time. In this section, we describe in detail our entire model compression method which is made of four parts, and each part corresponds to one facet of merits.
@@
Preferable compression ratio
The core objective of the algorithm is to identify the smallest $\theta^{*}$ for points in hypercubes $[0,1]^{K}$ under the condition that $\epsilon$ is sufficiently small and aiming for the highest feasible dimension $K$. To this end, we consider the distribution of weights of the original network by binning weights into three groups based on magnitudes of weight values.
  • First, we flatten all parameters into a one-dimensional vector $\mathbf{W}=[W_{1},W_{2},\cdots,W_{N}]$, where $N$ is the total number of weights, and then split it into groups $\left[\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}\right]$, where $W^{(g)}=[W_{K(g-1)+1},W_{K(g-1)+2},\cdots,W_{K(g-1)+K}],g=1,2,\cdots,G$ is of $\mathbb{R}^{1\times K}$, and $G$ is the number of groups. For any $\mathbf{W}^{(g)}$, there exists an integer $\theta^{*}_{g}$ such that $|\mathbf{W}^{(g)}-\tau(\theta^{*}_{g}\cdot\mathbf{a})|<\epsilon,g=1,\cdots,G$ and $\mathbf{a}=[a_{1},a_{2},\cdots,a_{K}]$. Finally, the vector $\left[\theta^{*}_{1},\theta^{*}_{2},\cdots,\theta^{*}_{G}\right]$ represents a compression to $\mathbf{W}$. Take $K=2$ and $\mathbf{W}=[W_{1},W_{2},W_{3},W_{4}]=[0.1,0.2,0.4,0.5]$ as an example, we split $\mathbf{W}$ as $[\mathbf{W}^{(1)},\mathbf{W}^{(2)}]$, where $\mathbf{W}^{(1)}=[W_{1},W_{2}]$ and $\mathbf{W}^{(2)}=[W_{3},W_{4}]$, and derive $\theta^{*}_{1}$ and $\theta^{*}_{2}$ for $\mathbf{W}^{(1)}$ and $\mathbf{W}^{(2)}$, respectively. As Figure 7 shows, we can find a $\theta$ for $(x,y)=(\tau(\theta/(\pi+1)),\tau(\theta/(\pi+2)))$ that approximates a point in 2D space.
@@
  • In our algorithm, we set $K=2$ and replace the unit square $[0,1]^{2}$ with a more flexible box $[a,b]\times[c,d]$. Specifically, given $\mathbf{W}$, $[a,b]\times[c,d]$ is defined as a square with side length $l$, centered at $(\bar{x_{i}},\bar{y_{i}})$, where $(\bar{x_{i}},\bar{y_{i}})$ is the centroid of two-dimensional points $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$. $l$ is a hyperparameter and typically set to 0.01 for optimal results. Consequently, the function $\tau(x):=x-\lfloor x\rfloor$ in the ergodic theorem is redefined as $\tau(z):=f(z)+([\bar{x_{i}},\bar{y_{i}}]-[\frac{l}{2},\frac{l}{2}])$, where $f(z)=z\bmod l$.
  • In the ergodic theorem, $\mathbf{a}=[a_{1},a_{2}]=\mathbf{d}\cdot I$ is a two-dimensional directional vector, where $\mathbf{d}$ and $I$ determine the direction vector and step size for the movement of curves defined by the parametric equation, respectively. Since $\theta$ is an integer no more than $U$, at most we have $U+1$ points in two-dimensional space $(\tau(a_{1}\theta),\tau(a_{2}\theta)),\theta=0,1,\cdots,U$. In our experiment, we define an optimal vector $\mathbf{a}$ based on $U$, such that the distribution of $U+1$ points is more uniform, thereby minimizing the maximum error as much as possible. Specifically,
@@
  • If we construct a square with its center at the centroid $(\bar{x_{i}},\bar{y_{i}})$ and side length $l$, it is highly probable that many points in $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$ will fall outside this square. Therefore, we need to first scale these points into the square using a scaling factor $s$ to compress, and then scale the substituted points $\tau(\theta^{*}_{g}\cdot\mathbf{a})$ back during decompression. While the error rate keeps intact, this inversion will amplify the error $\epsilon$, where $\epsilon=|W^{(g)}-\tau(\theta^{*}_{g}\cdot a)|$. It should be noted that, assuming $\mathbf{W}^{(F)}$ is the point farthest from the centroid $(\bar{x_{i}},\bar{y_{i}})$ in $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$, we define the farthest distance $l_{f}$ as $|\mathbf{W}^{(F)}-[\bar{x_{i}},\bar{y_{i}}]|$, and then the scaling factor $s=\frac{l}{2\cdot l_{f}}$. The specific formula derivation is as follows:
    $\mathbf{W}^{(in)}$ $=\mathbf{W}^{(in)}-C$ (5)
    $=C\mathbf{W}^{(out)}\cdot s$
    $=C\mathbf{W}^{(out)}\cdot\frac{l}{2\cdot l_{f}},$
    where $C$ is the centroid node, $\mathbf{W}^{(out)}$ is a node ouside the squre, $\mathbf{W}^{(in)}$ is the scaled node of $\mathbf{W}^{(out)}$.
    $\mathbf{W}^{(out)*}$ $=\frac{C\mathbf{W}^{(in)*}}{s}+C$ (6)
    $=\frac{\mathbf{W}^{(in)}+\epsilon}{s}+C$
    Therefore, we can derive the error:
    $=|\frac{\epsilon}{s}|=|\frac{2\cdot\epsilon\cdot l_{f}}{l}|$
    From (7), we can observe that the farthest distance $l_{f}$ determines the error of estimating $\mathbf{W}^{(out)}$ after scaling back.
    Therefore, as shown in Fig. 9 in our algorithm, we classify the points outside the square into $M$ categories based on their distances from the centroid $C$, where $M$ is a hyperparameter, so that the points in different categories can utilize different farthest distance $l_{fm}$ to define different scale factor $s_{m}$. This can mitigate the adverse impact of the global farthest distance $l_{f}$ on the final error as much as possible. This method is highly effective in practice, as a significant portion of the parameter nodes $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$ exhibit a pattern that most points are close to the centroid.
    Specifically, assuming the points outside the square are divided into $M$ categories and $W^{(m)}$ is a point that belongs to the $m$-th category, the scaling factor $s_{m}$ for this point is defined as
    $s_{m}=\frac{l/2}{l/2+(l_{f}/K)\cdot m}.$ (8)
    The final compression $\theta_{m}^{*}$ is given by the following formula:
    $\theta_{m}^{*}=m\cdot U+\lambda_{m}^{*},$ (9)
No fine-tuning
Our method does not require fine-tuning of the model, and can directly use the trained model for compression. This greatly simplifies the model compression process and avoids the time and computing resource consumption caused by retraining the model. The reasons why our algorithm requires no retraining are mainly because our algorithm can guarantee the error rate is low by optimizing the direction of $\mathbf{a}$, the side length $l$, and $U$. The error rate of points in different classes is around $(\frac{l}{U})/l=\frac{1}{U}$.
Affordable inference time
@@
At first glance, one may think that model compression based on hyperfunctions suffers from a slow inference time, as this technique needs to restore parameters before inference, which adds another level of computation. Here, we propose a scheme that leverages the intrinsic hierarchical structure of a network to greatly reduce the inference time. Our scheme parallelizes parameter decompression and inference. As shown in Figure 10, while the parameters of later layers (except the first part) are restored, the inference operation in earlier layers is also carried out simultaneously. As long as we can recover the parameters of the current layer before the inference arrives at the current layer, there is no waste of time. Thus, theoretically, the inference time of using our algorithm only increases quite moderately. The increment is the time used to restore the parameters of the first layer.
@@
Figure 10: The rapid growth of LLM sizes has outpaced the growth of GPU memory, creating challenges in serving these increasingly massive models.
@@
Based on the above solution, we need to shorten the decompression time. First, the entire decompression process is implemented using matrix operations, which can significantly speed up the computation. Second, the process of reading files from storage and preprocessing them is relatively time-consuming, and repeats each time the function is called. Therefore, we optimize this process, performing file reading and preprocessing operations only when the function is called for the first time, and storing the processing results in the cache. When the function is called again, the preprocessed intermediate results are directly captured from the cache, further shortening the inference time.
Short Compression Time
@@
  • As mentioned earlier, when encoding the parameter vector $\mathbf{W}$, we first convert it into many points $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$. Then, given $\mathbf{W}^{(g)}$, we search $u$ from $\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})$ to approximate $\mathbf{W}^{(g)}$. Next, instead of repeating searching for every point, we store $\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})$ in the format of k-d trees [17]. Thus, we can fast determine $\tau(u\cdot\mathbf{a})\in[\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})]$ that has the shortest distance to $\mathbf{W}^{(g)}$.
@@
    where $s_{g}\in F$ represents the scale factor of $\mathbf{W}^{(g)}$, $m_{g}$ means the category of node $\mathbf{W}^{(g)}$ and the points located within the center square are classified as category $m_{g}$.

### v2

@@ II. Limit of the Compression Ratio of Pruning Convolutional Networks
The class of graphs we investigate is $G_{r}=(V_{r},E_{r}),~{}r\in\mathbb{Z}_{\geq 2}$, where $V_{r}=\mathbb{Z}^{2},E_{r}=\{\langle(m,n),(m+1,n+i)\rangle\mid m,n\in\mathbb{Z},i\in{0,1,\cdots,r-1}\}$, that is, a convolution with the kernel size $r$ and stride 1. Figure 6 and 6 show the cases of $r=2$ and $r=3$.
@@
Earlier, researchers have mainly focused on planar graphs, but we observe that $G_{r}$ is a planar graph only when $r=2$, and is non-planar for $r\geq 3$, parts of $G_{2}$ and $G_{3}$ are shown in Figures 6 and 6. Therefore, many methods, such as using the duality of planar graphs and the star-triangular transition, cannot be applied. In the following, we will first solve the critical probability for the case of $r=2$, and then study the upper and lower bounds of the critical probability for $r\geq 3$. Our method is to degenerate a non-planar graph into a planar graph by removing edges.
@@
Remark. From the viewpoint of percolation theory, the compression ratio of pruning a randomly generated convolutional network is $1/p_{0}<\frac{1}{p_{c,r}}<\mu_{r}$, which is a fixed constant given $r$. Our analysis is based on an assumption that ignores the association between two edges. In reality, when an edge with a larger weight is pruned, edges with smaller weights are also pruned. Future work will be considering the association to have a more accurate estimation for the pruning bound.
@@ III. Algorithmic Details
Anchored in the hyperfunction driven by ergodic theory mentioned in the main body, we design a general-purpose model compression method that enjoys the following merits (PNAS): Preferable compression ratio, No post-hoc retraining, Affordable inference time, and Short compression time. In this section, we describe in detail our entire model compression algorithm which is made of four parts, and each part corresponds to one facet of merits.
@@
Compression via Ergodicity $\to$ Preferable compression ratio
The core objective of the algorithm is to identify the smallest $\theta^{*}$ for points in hypercubes $[0,1]^{K}$ under the condition that $\epsilon$ is sufficiently small and $K$ is as high as possible. To this end, we consider the distribution of weights of the original network by binning weights into different groups based on magnitudes of weight values.
  • First, we flatten all parameters into a one-dimensional vector $\mathbf{W}=[W_{1},W_{2},\cdots,W_{N}]$, where $N$ is the total number of weights, and then split it into groups $\left[\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}\right]$, where $W^{(g)}=[W_{K(g-1)+1},W_{K(g-1)+2},\cdots,W_{K(g-1)+K}],g=1,2,\cdots,G$ is of $\mathbb{R}^{1\times K}$, and $G$ is the number of groups. For any $\mathbf{W}^{(g)}$, there exists an integer $\theta^{*}_{g}$ such that $|\mathbf{W}^{(g)}-\tau(\theta^{*}_{g}\cdot\mathbf{a})|<\epsilon,g=1,\cdots,G$ and $\mathbf{a}=[a_{1},a_{2},\cdots,a_{K}]$. Finally, the vector $\left[\theta^{*}_{1},\theta^{*}_{2},\cdots,\theta^{*}_{G}\right]$ represents a compression to $\mathbf{W}$. Take $K=2$ and $\mathbf{W}=[W_{1},W_{2},W_{3},W_{4}]=[0.1,0.2,0.4,0.5]$ as an example, we split $\mathbf{W}$ as $[\mathbf{W}^{(1)},\mathbf{W}^{(2)}]$, where $\mathbf{W}^{(1)}=[W_{1},W_{2}]=[0.1,0.2]$ and $\mathbf{W}^{(2)}=[W_{3},W_{4}]=[0.4,0.5]$, and derive $\theta^{*}_{1}$ and $\theta^{*}_{2}$ for $\mathbf{W}^{(1)}$ and $\mathbf{W}^{(2)}$, respectively. As Figure 7 shows, we can find a $\theta$ such that $(\tau(\theta/(\pi+1)),\tau(\theta/(\pi+2)))$ can approximate any point in 2D space.
@@
  • In our algorithm, we set $K=2$ and replace the unit square $[0,1]^{2}$ with a more flexible box $[a,b]\times[c,d]$. Specifically, given $\mathbf{W}$, $[a,b]\times[c,d]$ is defined as a square with side length $l$, centered at $(\bar{x_{i}},\bar{y_{i}})$, where $(\bar{x_{i}},\bar{y_{i}})$ is the centroid of two-dimensional points $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$. $l$ is a hyperparameter typically set to 0.01 for optimal results. Consequently, the function $\tau(x):=x-\lfloor x\rfloor$ in the ergodic theorem is generalized as $\tau([x,y]):=f([x,y])+([\bar{x_{i}},\bar{y_{i}}]-[\frac{l}{2},\frac{l}{2}])$, where $f(x)=x\bmod l$.
  • If we construct a square with its center at the centroid $(\bar{x_{i}},\bar{y_{i}})$ and the side length $l$, it is highly probable that many points in $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$ will fall outside this square. Therefore, we need to first scale these points into the square using a scaling factor $s$ to compress, and then scale the substituted points $\tau(\theta^{*}_{g}\cdot\mathbf{a})$ back during decompression. While the error rate remains intact, this inversion will amplify the error $\epsilon$, where $\epsilon=|W^{(g)}-\tau(\theta^{*}_{g}\cdot\mathbf{a})|$. Assuming $\mathbf{W}^{(F)}$ is the point farthest from the centroid $(\bar{x_{i}},\bar{y_{i}})$ in $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$, we define the farthest distance $l_{f}$ as $|\mathbf{W}^{(F)}-[\bar{x_{i}},\bar{y_{i}}]|$, and then the scaling factor $s=\frac{l}{2\cdot l_{f}}$. The specific formula derivation is as follows:
    $\mathbf{W}^{(in)}$ $=\mathbf{W}^{(in)}-C$ (2)
    $=C\mathbf{W}^{(out)}\cdot s$
    $=C\mathbf{W}^{(out)}\cdot\frac{l}{2\cdot l_{f}},$
    where $C$ is the centroid node, $\mathbf{W}^{(out)}$ is a node ouside the squre, and $\mathbf{W}^{(in)}$ is the scaled node of $\mathbf{W}^{(out)}$.
    $\mathbf{W}^{(out)*}$ $=\frac{C\mathbf{W}^{(in)*}}{s}+C$ (3)
    $=\frac{\mathbf{W}^{(in)}+\epsilon}{s}+C.$
    Therefore, we can derive the error:
    $=|\frac{\epsilon}{s}|=|\frac{2\cdot\epsilon\cdot l_{f}}{l}|$
    From (4), we can observe that the farthest distance $l_{f}$ affects the error of estimating $\mathbf{W}^{(out)}$ after scaling back.
    Therefore, as shown in Fig. 8 in our algorithm, we classify the points outside the square into $M$ categories based on their distances from the centroid $C$, where $M$ is a hyperparameter, so that the points in different categories can utilize different farthest distance $l_{fm}$ to define different scale factor $s_{m}$. This is better than using the global farthest distance $l_{f}$ to control the compression error. This method is highly effective in practice, as a significant portion of the parameter nodes $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$ exhibit a pattern that most points are close to the centroid.
    Specifically, assuming that the points outside the square are divided into $M$ categories and $W^{(m)}$ is a point under the $m$-th category, the scaling factor $s_{m}$ for this point is defined as
    $s_{m}=\frac{l/2}{l/2+(l_{f}/K)\cdot m}.$ (5)
    The final compression $\theta_{m}^{*}$ is given by the following formula:
    $\theta_{m}^{*}=m\cdot U+\lambda_{m}^{*},$ (6)
Adjusting Irrational Direction $\to$ No fine-tuning
Our method does not require fine-tuning of the model, and can directly use the trained model for compression. This greatly simplifies the model compression process and avoids the time and computing resource consumption caused by retraining the model. One reason accounting for why our algorithm requires no retraining is mainly because our algorithm can guarantee the error rate is low by optimizing the direction of $\mathbf{a}$, the side length $l$, and $U$. The error rate of points in different classes is always no more than $(\frac{l}{U})/l=\frac{1}{U}$.
Specifically, in the ergodic theorem, $\mathbf{a}=[a_{1},a_{2}]=\mathbf{d}\cdot I$ is a two-dimensional directional vector, where $\mathbf{d}$ and $I$ determine the direction vector and step size for the movement of curves defined by the parametric equation, respectively. Since $\theta$ is an integer no more than $U$, at most we have $U+1$ points in two-dimensional space $(\tau(a_{1}\theta),\tau(a_{2}\theta)),\theta=0,1,\cdots,U$. In our experiment, we define an optimal vector $\mathbf{a}$ based on $U$, such that the distribution of $U+1$ points is more uniform, thereby minimizing the maximum error as much as possible. Specifically,
@@
Conducting Inference and Decompression Simultaneously $\to$ Affordable inference time
@@
At first glance, one may think that model compression based on hyperfunctions suffers from a slow inference time, as this technique needs to restore parameters before inference, which adds another level of computation. Here, we propose a scheme that leverages the intrinsic hierarchical structure of a network to greatly reduce the inference time. Our scheme parallelizes parameter decompression and inference. As shown in Figure 10, while the parameters of later layers (except the first layer) are restored, the inference operation in earlier layers is also carried out simultaneously. As long as we can recover the parameters of the current layer before the inference arrives at this layer, there is no waste of time. Thus, theoretically, the inference time of using our algorithm only increases quite moderately. The increment is the time used to restore the parameters of the first layer.
@@
Figure 10: The rapid growth of LLM’s size has outpaced the growth of GPU memory, creating challenges in serving these increasingly massive models.
@@
Based on the above solution, we need to shorten the decompression time. First, the entire decompression process is implemented using matrix operations, which can significantly speed up the computation. Second, the process of reading files from storage and preprocessing them is relatively time-consuming, and repeats each time the function is called. Therefore, we optimize this process, performing file reading and preprocessing operations only when the function is called for the first time, and storing the processing results in the cache. When the function is called again, the preprocessed intermediate results are directly captured from the cache, further accelerating the inference.
KD Tree + Parallelization $\to$ Short Compression Time
@@
  • As mentioned earlier, when encoding the parameter vector $\mathbf{W}$, we first convert it into many points $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$. Then, given $\mathbf{W}^{(g)}$, we search $u$ from $\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})$ to approximate $\mathbf{W}^{(g)}$. Next, instead of repeating searching for every point, we store $\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})$ in the format of k-d trees [19]. Thus, we can fast determine $\tau(u\cdot\mathbf{a})\in\{\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})\}$ that is closet to $\mathbf{W}^{(g)}$.
@@
    where $s_{g}\in F$ represents the scale factor of $\mathbf{W}^{(g)}$ and $m_{g}$ means the category of node $\mathbf{W}^{(g)}$. The points located within the center square are classified as category $m_{g}$.